# Syntax-Guided Procedural Synthesis of Molecules

### v1

@@ Abstract
Designing synthetically accessible molecules and recommending analogs to unsynthesizable molecules are important problems for accelerating molecular discovery. We reconceptualize both problems using ideas from program synthesis. Drawing inspiration from syntax-guided synthesis approaches, we decouple the syntactic skeleton from the semantics of a synthetic tree to create a bilevel framework for reasoning about the combinatorial space of synthesis pathways. Given a molecule we aim to generate analogs for, we iteratively refine its skeletal characteristics via Markov Chain Monte Carlo simulations over the space of syntactic skeletons. Given a black-box oracle to optimize, we formulate a joint design space over syntactic templates and molecular descriptors and introduce evolutionary algorithms that optimize both syntactic and semantic dimensions synergistically. Our key insight is that once the syntactic skeleton is set, we can amortize over the search complexity of deriving the program’s semantics by training policies to fully utilize the fixed horizon Markov Decision Process imposed by the syntactic template. We demonstrate performance advantages of our bilevel framework for synthesizable analog generation and synthesizable molecule design. Notably, our approach offers the user explicit control over the resources required to perform synthesis and biases the design space towards simpler solutions, making it particularly promising for autonomous synthesis platforms.
@@ 1 Introduction
The discovery of new molecular entities is central to advancements in fields such as pharmaceuticals [72, 40], materials science [26, 31], and environmental engineering [73, 67]. Traditional make-design-test workflows for molecular design typically rely on labor-intensive methods that involve a high degree of trial and error [52]. Systematic and data-efficient approaches that minimize costly experimental trials are the key to accelerating these processes [11, 12, 22]. In recent years, a large number of molecular generative models has been proposed [15, 41, 55, 68, 38, 51, 39, 32, 33, 25, 59]. However, few of their outputs are feasible to make and proceed to experimental testing due to their lack of consideration for synthesizability [20]. This has motivated methods that integrate design and synthesis into a single workflow, aiming to optimize both processes simultaneously [65, 27, 6, 2, 3, 21, 60] which significantly closes the gap between the design and make steps, reducing cycle time significantly [36, 66, 43]. These methods still face computational challenges, particularly in navigating the combinatorial explosion of potential synthetic pathways [56].
@@
Inspired by techniques in program synthesis, particularly Syntax-Guided Synthesis (SyGuS) [1], our method decouples the syntactical template of synthetic pathways (the skeleton) from their chemical semantics (the substance). This bifurcation allows for a more granular optimization process, wherein the syntactical and semantic aspects of reaction pathways can be optimized independently yet synergistically. Our methodology employs a bilevel optimization strategy, wherein the upper level optimizes the syntactic template of the synthetic pathway, and the lower level fine-tunes the molecular descriptors within that given structural framework. This dual-layered approach is facilitated by a surrogate policy, implemented via a graph neural network, that propagates embeddings top-down following the topological order of the syntactical skeleton. This ensures that each step in the synthetic pathway is optimized in context, respecting the overarching structural strategy while refining the molecular details. We address the combinatorial explosion in the number of programs using tailored strategies for fixed horizon Markov decision process (MDP) environments. This algorithm amortizes the complexity of the search space through predictive modeling and simulation of Markov Chain Monte Carlo (MCMC) processes [47, 28, 24], focusing on the generation and evaluation of syntactical skeletons. By leveraging the inductive biases from retrosynthetic analysis without resorting to retrosynthesis search, our approach combines accuracy and efficiency in “synthesizing" synthetic pathways. In summary, the contributions of this work are:
@@
  • We propose a bilevel framework that decouples the syntactical skeleton of a synthetic tree from the program semantics.
  • We introduce amortized algorithms within the bilevel framework for the tasks of synthesizable analog recommendation and synthesizable molecule design.
  • We demonstrate improvements across multiple dimensions of performance for both tasks.
  • We include in-depth visualizations and analyses for understanding the source of our method’s efficacy as well as its limitations.
@@ 2 Related Works
2.1 Synthesis Planning
Prior works model synthetic pathways using a discrete data structure known as the synthetic tree. The root node is the target molecule, leaf nodes are building blocks, and intermediate nodes can be either reactions or intermediates. This task is to infer a synthesis tree that reconstructs a target molecule $M\in\mathcal{M}$, where $\mathcal{M}$ is the space of all molecules (estimated to be $10^{30-100}$). Retrosynthetic analysis is a branch of organic chemistry that centers around designing synthesis pathways for a target molecule through backward reasoning steps [13]. Computer-assisted retrosynthetic analysis has developed through the decades [14] in tandem with computers, and is now known as retrosynthetic planning due to its resemblance to more classical tests of AI based around planning. State-of-the-art retrosynthesis planning algorithms today use neural networks to learn complex transformation patterns over the vast space of molecular structures, and the field has gained attention in recent years [9] as machine learning has begun to transform drug discovery and materials design.
@@ 2.2 Synthesizable Analog Generation
The problem of synthesizable analog generation aims to find molecules close to the target molecule which are synthesizable. Note that the constraint of synthesizable delinates this problem from conditional molecule generation, but works such as [50] attempt to bridge these two. As retrosynthetic planning is done by working backwards (top-down), partial success is not straightforward to define. In other domains, procedural modeling is a bottom-up generation process that generates analogs by design [45, 46, 48, 44]. Thus, synthesizable analog generation has warranted more specialized methods. Prior works such as [16, 37] address this by starting from an existing retrosynthesis route, and performing alteration of the route. This constrained approach limits the diversity of analogs severely. Instead, we neither start from a search route nor constrain the search route, but instead extract analogs via iterative refinement of the program’s syntactical skeleton with inner-loop decoding of the program semantics in a bilevel setup. Our framework simultaneously handles analog generation and, as a special case, synthesis planning. We implement the iterative refinement phase using a MCMC sampler with a stationary distribution governed by similarity to the target being conditioned on. This is a common technique used to search over procedural models of buildings, shapes, furniture arrangements, etc., [46, 61, 69] and we showcase its efficacy for the new application domain of molecules.
@@ 2.3 Synthesizable Molecule Design
The problem of synthesizable molecule design is to optimize for the synthetic pathway which produces a molecule that maximizes some property oracle function. Note that unlike generic molecular optimization approaches, the design space is reformulated to guarantee synthesizability by construction. The early works to follow this formulation [65, 27, 6] use machine learning to assemble molecules by iteratively selecting building blocks and virtual reaction templates to enumerate a library, with recent works such as [60] obtaining experimental validation. The key computational challenge these methods must address is how to best navigate the combinatorial search space of synthetic pathways. Prior works that do bottom-up generation of synthetic trees [2, 3] probabilistically model a synthetic tree as a sequence of actions. These works adopt an encoder-decoder approach to map to and from a latent space, e.g., using a VAE, assuming a smooth mapping between continuous latent space to a complex and highly discontinuous combinatorial space. This results in low reconstruction accuracy, hindering the method on the task of conditional generation. SynNet [21] instead formulates the problem as an infinite-horizon MDP and do amortized tree generation conditioned on a Morgan fingerprint. This enables a common framework for solving both tasks. However, we show improvements on both analog generation and synthesizable molecule design in terms of reconstruction accuracy and diversity through our novel formulation.
@@ 2.4 Program Synthesis
Program synthesis is the problem of synthesizing a function $f$ from a set of primitives and operators to meet some correctness specification. A program synthesis problem entails: a background theory $\mathsf{T}$ which is the vocabulary for constructing formulas, a correctness specification: a logical formula involving the output of $f$ and $\mathsf{T}$, and a set of expressions $L$ that $f$ can take on described by a context-free grammar $G_{L}$. In molecular synthesis, we can formulate $\mathsf{T}$ as containing operators for chemical reactions, constants for reagents, molecular graph isomorphism checking comparisons, etc. The correctness specification for finding a synthesis route for $M$ is simply $f(\{B\})=M$ (where $\{B\}$ is a set of building blocks) and we seek to find an implementation $f$ from $L$ to meet the specification. A related but coarser specification is to match the fingerprint $X$ of some molecule: $FP(f(\{B\}))=X$, and as shown by [21], this relaxed formulation enables both analog generation and GA-based molecule optimization. Our key innovation takes inspiration from the line of work surrounding syntax-guided synthesis [1, 53] (SyGuS). Syntax guidance explicitly constrains $G_{L}$ which reduces the set of implementations $f$ can take on [1], enabling more accurate amortized inference. Further discussion on the connections between program synthesis and molecular synthesis is in App C.
Figure 1: We introduce program synthesis terminology for modeling synthesis pathways.
@@ 3.1 Problem Definition
Synthesis Planning
This task is to infer the program as well as the inputs that reconstructs a target molecule. When the outcome is binary (whether a certificate to exactly reconstruct the target is found), this problem is known as synthesis planning. When the outcome can be between 0 and 1, e.g., similarity to target molecule, this problem is known as synthesizable analog generation.
Synthesizable Molecule Design
This task is to optimize for synthesizable molecules that maximize a property oracle function. This is a constrained optimization problem where the design space ensures synthesizability by construction.
@@ 3.2 Solution Overview
In our work, we use expert-defined reaction templates, a popular abstraction codifying deterministic graph transformation patterns in the SMIRKS formal language. SMIRKS assumes the reactants are ordered (for defining how atoms and bonds are transformed from reactants to products). Since templates are designed to handle the most common and most robust chemical transformations, ours are restricted to uni-molecular (e.g., isomerization, ring closure, or fragmentation reactions) or bi-molecular (e.g., additions, substitutions, or coupling reactions) reactions. Denoting the set of reactions as $\mathcal{R}$ and the set of building blocks as $\mathcal{B}$, we obtain a compact yet expressive design space $\mathcal{P}$: all non-trivial, attributed binary trees where each node corresponds to a reaction. The problem of synthesis planning is to infer the program and appropriate inputs $\mathcal{F}\colon\mathcal{M}\rightarrow\mathcal{P}\times 2^{\mathcal{B}}$, i.e., $(P,[b_{i}\in\mathcal{B}])$ such that $B$ can be assigned to the leaf nodes of $P$ and running the reaction(s) in $P$ in a bottom-up manner (by recursively feeding the product of a node’s reaction to its parent) produces $\hat{M}$ which minimizes $\text{dist}(\hat{M},M)$. We call the space of $\mathcal{P}$ program space to motivate our solution by drawing a parallel to program synthesis literature. Moving forwards, we define exec over the image of $\mathcal{F}$ as the output of executing the program $P$ on the inputs $B$. Equivalently, we use the functional notation $P(B)$. The basic terminology is summarized in Figure 1. Our featurization of $\mathcal{M}$ is chosen to be $\mathcal{X}$ the set of all Morgan fingerprints with a fixed radius and number of bits. This is a common representation of molecules for both predictive tasks and design tasks. Our parameterized surrogate procedure also takes $\mathcal{X}$ as its domain, i.e., $F_{\Theta,\Phi,\Omega}\colon\mathcal{X}\rightarrow\mathcal{P}\times 2^{\mathcal{B}}$, for both tasks.
@@ 3.3 Bilevel Syntax-Semantic Synthesizable Analog Generation
We propose a multi-step factorization to synthesize a program given a molecule $M\in\mathcal{M}$ by first choosing the syntactical skeleton of the solution, then filling in the specific operations and literals. Following the terminology of Section 2.4, we aim to parameterize the derivation procedure of the context-free grammar $G_{\mathcal{P}}$. We define the grammar $G_{\mathcal{P}}$ explicitly and discuss how our syntax-guided approach constrains the grammar in App D. In the following sections, we introduce the main ideas for how we parameterize the derivation of the syntax-guided grammar when conditioned on an input molecule $M$.
The first step $Q\colon\mathcal{M}\rightarrow\mathcal{T}$, where $\mathcal{T}$ is the space of all non-trivial binary trees, is a policy to choose which production rule to apply first. In other words, it recognizes the most likely syntax tree of $P$, given an input $M$. We parameterize $Q_{\Theta}\colon\mathcal{M}\rightarrow\mathcal{T}_{k}$ with a standard MLP for classification. When doing bilevel optimization, this first step is skipped as a sample $T$ is given. The second step is to use $G_{R}\colon\mathcal{T}\times\mathcal{M}\rightarrow\mathcal{P}$ to fill in the reactions. The final step, $G_{B}\colon\mathcal{M}\times\mathcal{P}\rightarrow(P,B)$ produces the final AST, where leaf nodes are literals (building blocks) and non-leaf nodes are operators (reactions). When performing inner-loop only, $F(M)\coloneqq(P_{M;Q},G_{B}(M,P_{M;Q}))$ where $P_{M;Q}\coloneqq G_{R}(Q(M),M)$. However, $Q(M)$ is not necessarily well-defined: there can be multiple skeletons $\{T(;M)\}$ per molecule $M$ (see App B for further investigation). Instead of direct prediction, our outer loop enables systematic exploration over $\mathcal{T}$ via invocation of the inner-loop in a bilevel setup. In this setup, we decouple $T$ from $F$, i.e., $F(M,T)\coloneqq(P_{M},G_{B}(M,P_{M}))$ where $P_{M}\coloneqq G_{R}(T,M)$.
3.3.1 Outer-Loop: Syntax Tree Structure Optimization

### v2

# Procedural Synthesis of Synthesizable Molecules

@@ Abstract
Designing synthetically accessible molecules and recommending analogs to unsynthesizable molecules are important problems for accelerating molecular discovery. We reconceptualize both problems using ideas from program synthesis. Drawing inspiration from syntax-guided synthesis approaches, we decouple the syntactic skeleton from the semantics of a synthetic tree to create a bi-level framework for reasoning about the combinatorial space of synthesis pathways. Given a molecule we aim to generate analogs for, we iteratively refine its skeletal characteristics via Markov Chain Monte Carlo simulations over the space of syntactic skeletons. Given a black-box oracle to optimize, we formulate a joint design space over syntactic templates and molecular descriptors and introduce evolutionary algorithms that optimize both syntactic and semantic dimensions synergistically. Our key insight is that once the syntactic skeleton is set, we can amortize over the search complexity of deriving the program’s semantics by training policies to fully utilize the fixed horizon Markov decision process imposed by the syntactic template. We demonstrate performance advantages of our bi-level framework for synthesizable analog generation and synthesizable molecule design. Notably, our approach offers the user explicit control over the resources required to perform synthesis and biases the design space towards simpler solutions, which is particularly promising for autonomous synthesis platforms. Code is at https://github.com/shiningsunnyday/SynthesisNet.
@@ 1 Introduction
The discovery of new molecular entities is central to advancements in fields such as pharmaceuticals (Zhavoronkov et al., 2019; Lyu et al., 2019), materials science (Hachmann et al., 2011; Janet et al., 2020), and environmental engineering (Zimmerman et al., 2020; Yao et al., 2021). Traditional make-design-test workflows for molecular design typically rely on labor-intensive methods that involve a high degree of trial and error (Sanchez-Lengeling & Aspuru-Guzik, 2018). Systematic and data-efficient approaches that minimize costly experimental trials are the key to accelerating these processes (Coley et al., 2020a; b; Gao et al., 2022). In recent years, a large number of molecular generative models has been proposed (De Cao & Kipf, 2018; Ma et al., 2018; Simonovsky & Komodakis, 2018; You et al., 2018; Li et al., 2018; Samanta et al., 2020; Liu et al., 2018; Jin et al., 2018; 2020; Guo et al., 2022; Sun et al., 2024). However, few of their outputs are feasible to make and proceed to experimental testing due to their lack of consideration for synthesizability (Gao & Coley, 2020). This has motivated recent methods that integrate design and synthesis into a single workflow, aiming to optimize both processes simultaneously (Button et al., 2019; Bradshaw et al., 2019; 2020; Gao et al., 2021; Swanson et al., 2024) which significantly closes the gap between the design and make steps, reducing cycle time significantly (Koscher et al., 2023; Volkamer et al., 2023; McCorkindale, 2023). This development is spurred by the curation of a small but robust collection of expert reaction templates that are inspired by real-world reactions and defined in close collaboration with chemists Hartenfeller et al. (2011). This workflow facilitates de novo applications by imposing synthetic accessibility by design, recasting molecular design as navigating the space of possible synthetic procedures over a set of building blocks and forward reaction steps, as defined in Vinkers et al. (2003). However, these methods still face computational challenges, particularly in navigating the combinatorial explosion of potential synthetic procedures (Smith, 1997).
@@
Inspired by techniques in program synthesis, particularly syntax-guided synthesis (Alur et al., 2013), our method decouples the syntactical template of a synthetic procedure (the skeleton) from their chemical semantics (the substance). This bifurcation allows for a more granular optimization process, wherein the syntactical and semantic aspects of reaction pathways can be optimized independently yet synergistically. Our methodology employs a bi-level optimization strategy. The upper level optimizes the syntactic template of the synthetic pathway, and the lower level fine-tunes the molecular descriptors within that given structural framework. This dual-layered approach is facilitated by a surrogate policy, implemented by a graph neural network, that propagates embeddings top-down following the topological order of the syntactical skeleton. This ensures that each step in the synthetic pathway is optimized in context, respecting the overarching structural strategy while refining the molecular details. We address the combinatorial explosion in the number of programs using tailored strategies for fixed horizon Markov decision process (MDP) environments. This algorithm amortizes the complexity of the search space through predictive modeling and simulation of Markov Chain Monte Carlo (MCMC) processes (Metropolis et al., 1953; Hastings, 1970; Gilks et al., 1995), focusing on the generation and evaluation of syntactical skeletons. By leveraging the inductive biases from retrosynthetic analysis without resorting to retrosynthesis search, our approach combines accuracy and efficiency in “synthesizing" synthetic pathways. In summary, the contributions of this work are:
  1. 1.
@@
  2. 2.
    We propose a bi-level framework that decouples the syntactical skeleton of a synthetic tree from the program semantics. Then, we introduce amortized algorithms within our framework for the tasks of synthesizable analog recommendation and synthesizable molecule design.
  3. 3.
    We demonstrate improvements across multiple dimensions of performance for both tasks, and include in-depth visualizations and analyses for understanding the source of our method’s efficacy as well as its limitations.
@@ 2.1 Synthesizable Analog Generation
The problem of synthesizable analog generation aims to find molecules close to the target molecule that are synthesizable. Closely related but distinct from this problem is computer-assisted retrosynthetic analysis, which has developed through the decades (Corey et al., 1985) in tandem with computers, and is now known as retrosynthetic planning due to its resemblance to more classical tests of AI based around planning. As retrosynthetic planning is done by working backwards (top-down), partial success is not straightforward to define. In other domains, procedural modeling is a bottom-up generation process that generates analogs by design (Merrell & Manocha, 2010; Merrell et al., 2011; Müller et al., 2006; Merrell, 2023). Thus, synthesizable analog generation warrants more specialized methods. Prior works such as Dolfus et al. (2022); Levin et al. (2023) address this by performing alterations of existing retrosynthesis routes, but this constrained approach severely limits the diversity of analogs. Instead, we neither start from a search route nor constrain the search route, but instead, extract analogs via the iterative refinement of the program’s syntactical skeleton with inner loop decoding of the program semantics in a bi-level setup. We implement the iterative refinement phase using an MCMC sampler with a stationary distribution governed by similarity to the target being conditioned on. This is a common technique used to search over procedural models of buildings, shapes, and furniture arrangements (Merrell et al., 2011; Talton et al., 2011; Yu et al., 2011), and we showcase its efficacy for the new application domain of molecules.
@@ 2.2 Synthesizable Molecule Design
The problem of synthesizable molecule design is to design the synthetic pathway, or program, whose output molecule optimizes some property oracle function. Note that unlike generic molecular optimization approaches, the design space is reformulated to guarantee synthesizability by construction. The early works to follow this formulation (Vinkers et al., 2003; Hartenfeller et al., 2011; Button et al., 2019) use machine learning to assemble molecules by iteratively selecting building blocks and virtual reaction templates to enumerate a library, with recent works such as Swanson et al. (2024) obtaining experimental validation. The key computational challenge these methods must address is how to best navigate the combinatorial search space of synthetic pathways. Prior efforts using bottom-up generation (Bradshaw et al., 2019; 2020) probabilistically model synthetic trees as sequences of actions. These adopt an encoder-decoder approach to map between a continuous latent space and a complex combinatorial space. This results in low reconstruction accuracy and hinders the method on the task of conditional generation. Instead, SynNet (Gao et al., 2021) admits a unified framework for solving analog generation and molecule design by formulating the problem as an infinite-horizon MDP and doing amortized tree generation conditioned on Morgan fingerprints. We show improvements to both tasks through our novel formulation.
@@ 2.3 Program Synthesis
Program synthesis is the problem of synthesizing a function $f$ from a set of primitives and operators to meet some correctness specification. A program synthesis problem entails: (1) a background theory $\mathsf{T}$ that is the vocabulary for constructing formulas, (2) a correctness specification, i.e., a logical formula involving the output of $f$ and $\mathsf{T}$, and (3) a set of expressions $L$ that $f$ can take on, described by a context-free grammar $G_{L}$. In molecular synthesis, we can formulate $\mathsf{T}$ as containing operators for chemical reactions, constants for reagents, molecular graph isomorphism checking, and so forth. The correctness specification for finding a synthesis route for a molecule $M$ is simply $f({\mathcal{B}})=M$, where ${\mathcal{B}}$ is a set of building blocks, and we seek to find an implementation $f$ from $L$ to meet the specification. A coarser specification is to match the molecule’s fingerprint, $\text{FP}(f({\mathcal{B}}))=\text{FP}(M)$, and as shown by Gao et al. (2021), this relaxed formulation enables both analog generation and fingerprint-based molecule optimization. Our key innovation takes inspiration from the line of work surrounding syntax-guided synthesis (Alur et al., 2013; Schkufza et al., 2013). Syntax guidance explicitly constrains $G_{L}$, which reduces the set of implementations $f$ can take on (Alur et al., 2013) and facilitates more accurate amortized inference. Further discussion on the connections between program synthesis and molecular synthesis and the similarities to the literature of retrosynthesis are given in App. D.
Figure 1: Program synthesis terminology for modeling synthesis pathways.
@@ 3.1 Problem Definition
For clarity, we recapitulate the problems of interest discussed in Section 2:
Synthesizable analog generation is the inverse task of inferring the program and inputs that best reconstructs a target molecule. Denoting the set of reactions as $\mathcal{R}$ and the set of building blocks as ${\mathcal{B}}$, we obtain a compact yet expressive design space ${\mathcal{P}}$: all non-trivial, attributed binary trees where each internal node corresponds to a reaction. Drawing parallels to the program synthesis literature, we call ${\mathcal{P}}$ the program space. The problem can now be formalized: given a space of molecules $\mathcal{M}$, learn a mapping $F\colon\mathcal{M}\rightarrow{\mathcal{P}}\times{\mathcal{B}}^{*},M\mapsto(P,B)$ such that $B$ can be assigned to the leaf nodes of $P$ and running the reaction(s) in $P$ in a bottom-up manner (by recursively feeding the product of a node’s reaction to its parent) produces a molecule $M_{*}$ with minimal “distance" to $M$. Using program execution notation, the objective is stated as: $\operatorname*{arg\,min}_{(P,B)\in{\mathcal{P}}\times{\mathcal{B}}^{*}}\text{dist}(P(B),M)$.
Synthesizable molecule design is the forward task of finding the program and inputs whose output optimizes a property oracle function. The oracle can represent property predictors, simulation, experimental validation, etc. but are black-box in nature. The objective is $\operatorname*{arg\,min}_{(P,B)}\text{Oracle}(P(B))$.
Moving forwards, we identify $(P,B)$ with its output $M_{*}$ when it simplifies notation. Stripping the semantics from a program leaves behind a syntactic skeleton, which lies in the space ${\mathcal{T}}$ of non-trivial binary trees. Figure 1 illustrates the discussed terminologies [^1] Please refer to Figure 3 of Gao et al. (2021) for example chemical illustrations of the two tasks..
@@ 3.2 Solution Overview
Herein, we use expert-defined reaction templates, a popular abstraction codifying deterministic graph transformation patterns in the SMIRKS formal language. SMIRKS assumes the reactants are ordered for defining how atoms and bonds are transformed from reactants to products. Since templates are designed to encompass the most common and robust chemical transformations, ours are restricted to uni-molecular (e.g., isomerization, ring closure, or fragmentation reactions) or bi-molecular (e.g., additions, substitutions, or coupling reactions) reactions. In practice, we featurize molecules using Morgan fingerprints with radius 2 and $d=2048$ bits, which is a common molecular representation in both predictive and design tasks. This means that $F$ is now technically a map over fingerprint space ${\mathcal{X}}\subseteq\{0,1\}^{d}$. It is then natural to use the Tanimoto distance between fingerprints as our notion of molecular distance.
@@ 3.3 Bilevel Syntax-Semantic Synthesizable Analog Generation
Given a molecule $M\in\mathcal{M}$, we aim to find a program and inputs $(P,B)$ whose output $M_{*}$ is most similar to $M$. Suppose first that the syntax $T$ of $P$ was given. Then, we would use two learned policies $\pi_{\mathcal{R}}$ and $\pi_{\mathcal{{B}}}$ that iteratively attribute reactions and building blocks to the nodes of $T$ in topological order (Section 3.3.2). Recalling Section 2.3, this can be seen as parameterizing the derivation process of a context-free grammar $G_{\mathcal{P}}$. We give further details in App. E and discuss how these derivations are constrained through syntax.
The application of our policies essentially specifies a syntax-conditioned program synthesis map $F\colon\mathcal{M}\times{\mathcal{T}}\to{\mathcal{P}}\times{\mathcal{B}}^{*}$. However, $T$ is not known during inference, so to remove this dependence, we consider two further strategies. One is to learn a classifier $\tau{}\colon\mathcal{M}\rightarrow{\mathcal{T}}$ to predict the most likely syntax tree of the program that produces $M$. We implement $\tau{}$ with a multi-layer perceptron (MLP) trained under a standard classification task. [^2] In practice, we restrict to a finite subset of ${\mathcal{T}}$ by imposing a maximum number of reaction nodes. However, a single prediction may be inadequate, since there can be multiple skeletons corresponding to the same molecule. Instead, our second strategy uses a bi-level setup, where the outer loop explores syntax space ${\mathcal{T}}$ through invocations of the inner loop $F$, which explores the program’s semantics. This approach is further made efficient by amortizing the training of the inner loop.
3.3.1 Outer loop: Syntax Tree Structure Optimization